OpenStudy (anonymous):

Verify that x-3 is a factor of 2x^3+x^2-x-3

OpenStudy (lalaly):

substitute x=3 in equation if its 0 then x-3 is a factor

OpenStudy (anonymous):

2*3^3+3^2-3-3=54+9-6=57, not a factor

OpenStudy (anonymous):

2*27+9*2-3-3=54+9-6=57 so it could'n solve in x=3 in equation then how is can u say it x-3 is a factor